Can someone show me how to do this...please show your work :)
1/4x+1/4y=-1
x-y=8
A)(-6,2)
B)(4,-4)
C)(2,-6)
D)(6,-2)
and
solve for R:
T=rs+s
A)r=T-1
B)r=T-s/s
C)r=s-T/s
D)r=s-T/T
thanks>>>

Answer by vidhyak(80)

1/4x+1/4y=-1
x-y=8
can be written as
x + y = -4 ...........equation 1
x - y = 8 ...........equation 2

Add eq1 and eq2
2x = 4
x = 2
Substitute x = 2 in either eq1 or eq 2

x + y = -4
2 + y = -4
y = -4 -2
y = -6

x= 2 y =-6 Therefore answer is C


solve for R:
T=rs+s
A)r=T-1
B)r=T-s/s
C)r=s-T/s
D)r=s-T/T
T = rs + s
rewriting the equation
rs + s = T
rs = T - s
r = (T-s)/s

Answer is B